# Towards Locally Deployable Fine-Tuned Causal Large Language Models for Mode Choice Behaviour

### v1

@@ 2.1 Experiment Design
To evaluate the effect of instruction strategy on the predictive performance of large language models (LLMs), two distinct prompting styles were implemented. The first adopts a hybrid structure inspired by Chain-of-Thought (CoT) and the Reason+Act (ReACT) prompting, wherein the model is encouraged to simulate a deliberative reasoning process before arriving at a decision. In this format, the prompt explicitly instructs the LLM to consider relevant trade-offs and contextual factors—such as time, cost, and purpose—based on the travel scenario and commuter profile, and to articulate a rationale before stating its selected mode. This approach facilitates cognitive transparency by allowing the model to express its internal reasoning prior to outputting a final decision.
@@
The second prompting style adopts a direct decision format, wherein the LLM is presented with the full scenario description including commuter attributes and available alternatives, and is asked to output only the selected mode, without generating intermediate reasoning. This method prioritizes succinctness and emulates standard classification behaviour, allowing for rapid assessments of model preference without the cognitive load of justifying choices.
@@
In addition to prompt style, the temperature parameter of the LLM was varied to explore the influence of generative randomness on prediction outcomes. In this context, temperature controls the stochasticity of the model’s token sampling process, with lower values (e.g., 0.5) promoting more deterministic and focused outputs, and higher values (e.g., 1.0) allowing for greater variation and exploratory generation. Temperature 0.50.50.5 was selected to assess performance under conservative, reproducible conditions, while temperature 1.01.01.0 was included to evaluate whether increased generative diversity may enhance generalization, particularly in cases involving ambiguity or competing trade-offs in the mode choice scenario.
@@
[…]
@@
[…]
Each experiment was subsequently evaluated by comparing the model generated predictions against the actual survey respondents’ mode choices. This evaluation framework allows for a structured and systematic assessment of each LLM’s predictive performance under varying learning strategies, prompting styles, and temperature settings. This allows the investigation into the behavioural fidelity of LLMs in emulating human travel decision making and highlights the conditions under which these models can offer the most accurate and generalizable representations of mode choice behaviour. The details of the evaluation framework are presented in Section 2.3
@@ 2. Targeted Similarity-Based Sampling
[…]
@@
Figure 2 illustrates the full local inference setup used in this study. It presents the pipeline from structured survey data ingestion to LLM querying and storage of predictions. The figure complements the formal equations and clarifies the logical and computational flow between components of the inference system.
Figure 2: Overall Experiment Design
@@ 2.3.1 LLMs Predictive Performance Evaluation
Let $\mathcal{D}_{\text{test}}=\{(d_{i},y_{i})\}_{i=1}^{N}$ be the test dataset, where $d_{i}$ is the input data (trip and sociodemographic attributes), and $y_{i}\in\{1,2,3\}$ denotes the ground-truth mode label and $C$ denote the total number of transport mode categories used in the classification task $c\in\{1,2,\dots,C\}$. Each evaluated model $\mathcal{M}^{(m)}$ generates a predicted mode $\hat{y}_{i}^{(m)}$ for each instance $i$. We compute the following standard classification metrics:
  • Accuracy:
@@
  • Recall (Macro):
@@
  • F1 Score (Macro):
[…]
  • F1 Score (Weighted): weighted by class prevalence in the ground truth.
In addition to these instance-level metrics, we compute distribution-level divergence between the predicted and true mode shares. Let:
  • $p_{c}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(y_{i}=c)$: empirical ground-truth distribution
  • $\hat{p}_{c}^{(m)}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(\hat{y}_{i}^{(m)}=c)$: predicted distribution by model $\mathcal{M}^{(m)}$
To handle zero-count issues, we apply Laplace smoothing with a small constant $\varepsilon=10^{-9}$:
$p_{c}^{\text{smooth}}=\frac{p_{c}+\varepsilon}{1+3\varepsilon},\quad\hat{p}_{c}^{(m),\text{smooth}}=\frac{\hat{p}_{c}^{(m)}+\varepsilon}{1+3\varepsilon}$ (18)
We then compute:
  • Mean Absolute Error (DistMAE):
    $\text{DistMAE}^{(m)}=\frac{1}{C}\sum_{c=1}^{C}\left|p_{c}-\hat{p}_{c}^{(m)}\right|$ (19)
  • Jensen-Shannon Divergence (JSD):
    $\text{JSD}^{(m)}=\frac{1}{2}\left[\sum_{c=1}^{C}p_{c}^{\text{smooth}}\log\left(\frac{p_{c}^{\text{smooth}}}{M_{c}}\right)+\hat{p}_{c}^{(m),\text{smooth}}\log\left(\frac{\hat{p}_{c}^{(m),\text{smooth}}}{M_{c}}\right)\right]$ (20)
    where $M_{c}=\frac{1}{2}(p_{c}^{\text{smooth}}+\hat{p}_{c}^{(m),\text{smooth}})$
  • Cross-Entropy:
    $\text{CE}^{(m)}=-\sum_{c=1}^{C}p_{c}^{\text{smooth}}\log\left(\hat{p}_{c}^{(m),\text{smooth}}\right)$ (21)
Each metric in the evaluation framework was selected for its ability to capture distinct yet complementary aspects of model performance. Accuracy provides an overall correctness measure but can be biased by class imbalance. Precision and recall help evaluate model performance on individual classes, with F1_Macro offering a class-averaged harmonic mean and F1_Weighted accounting for class imbalance by weighting each class’s F1 score by its support. On the distribution level, MAE quantifies the average deviation between predicted and true mode share proportions, while cross-entropy penalizes probabilistic mismatch and is sensitive to over or under-confidence in predictions. JSD, a symmetrized and smoothed variant of Kullback-Leibler divergence, offers a bounded and interpretable measure of similarity between the predicted and actual distributions. Together, these metrics ensure robust evaluation across both individual and aggregate behavioural patterns.
@@ 2.3.2 LLMs Reasoning Performance Evaluation
To complement the evaluation of predicted travel modes, we conducted a detailed reasoning analysis to assess the structure, diversity, and strength of the natural language justifications generated by each model. This analysis consisted of three components: (i) computation of an explanation strength index, (ii) topic modelling using BERTopic, and (iii) comparative visualizations across model configurations.
[…]
@@
Semantic Embedding and UMAP Projection: Each reasoning $z_{i}$ was encoded into a high-dimensional semantic vector $\mathbf{e}_{i}\in\mathbb{R}^{d}$ using the all-MiniLM-L6-v2 sentence transformer model. These embeddings were reduced to two dimensions via Uniform Manifold Approximation and Projection (UMAP) for visualization:
@@
[…]
Topic Modeling with BERTopic: The BERTopic algorithm was then applied using the sentence embeddings and UMAP-reduced vectors. This model clustered the explanations into coherent topics $T=\{t_{1},t_{2},\dots,t_{K}\}$, with each instance assigned a topic label $t_{i}\in T$. The number of topics was fixed to $K=8$. For each topic, a ranked list of representative keywords was extracted using class-based TF-IDF, enhancing interpretability.
Comparative Visualization and Analysis: To examine differences in explanation structure and content across models and prompting setups, we generated three visual analyses:
  • Topic Prevalence by Model: For each model $m$, we computed the proportion $\pi_{m,k}$ of reasonings assigned to topic $t_{k}$, facilitating comparison of dominant narrative patterns.
  • ESI Heatmap: Mean ESI values were aggregated across prompt types and temperature settings to assess how model configuration affects explanation strength.
  • UMAP Topic Scatter: The reduced embeddings $\mathbf{u}_{i}$ were visualized and color-coded by their assigned topic $t_{i}$, offering insight into semantic clustering.
Together, these components enabled a structured evaluation of the interpretive dimension of model outputs, allowing us to assess not only prediction accuracy but also the content richness and diversity of the generated rationales.
@@ 2.4 Fine-tunning
To adapt a foundational Large Language Mode(LLM) for the specialized task of travel mode choice prediction, a comprehensive finetuning methodology was implemented. This process involves the careful selection of a parameter-efficient tuning technique, a tailored training objective to prevent common issues with decoder-only models, and a robust evaluation framework to ensure model generalization.
The foundation for this research is the decoder-only transformer LLMs AKA Casual LLMs. Its extensive pre-training provides a rich basis of general knowledge, which we specialize through finetuning.
The training corpus consists of several thousand structured examples derived from travel survey responses. Each example is composed of: 1) an instruction, a detailed text prompt describing a traveler’s sociodemographic profile and a specific travel scenario with the transport options and their attributes (e.g., travel time, cost, frequency); and 2) a selected_mode, the ground-truth label corresponding to the mode choice.
For robust model development, the dataset was partitioned into a training set (90%) and a validation set (10%). This allows for performance monitoring on unseen data to prevent overfitting and to select the best-performing model checkpoint. This is separate from the 200 test set that was set a side for the finetuned model validation and comparison similar to the process described in section
@@ 2.4.1 Parameter Efficient Finetuning (PEFT) Strategy: QLoRA
Our parameter-efficient finetuning strategy is based on Quantized Low-Rank Adaptation (QLoRA) (Dettmers et al., 2023), an advanced method that enhances the original Low-Rank Adaptation (LoRA) framework (Hu et al., 2022) by integrating 4-bit quantization of the base model’s weights, thereby significantly reducing memory requirements. This strategy makes finetuning large models feasible on limited hardware (a single GPU with 12GB VRAM) by drastically reducing the number of trainable parameters.
@@
For this study, we configured LoRA with a rank $r=32$ and a scaling factor $\alpha=64$. These trainable matrices were applied to all major linear layers of the model architecture, including query, key, value, and output projections, as well as the feed forward network layers.
@@ 2.4.2 Training Objective and Label Leakage Mitigation
As the model was finetuned using a Causal Language Modelling (CLM) objective. A significant challenge when applying this objective to structured prediction tasks with decoder-only architectures is label leakage. This occurs when the ground-truth answer is exposed to the model as part of its input context, allowing the model to achieve high performance by simply copying the answer rather than learning the task.
[…]
@@
where $\theta$ represents the model’s trainable parameters, $P(\tau_{j}\mid\tau_{<j};\theta)$ is the predicted probability of token $\tau_{j}$, and $\mathbb{I}(\cdot)$ is an indicator function that equals 1 if the condition is true and 0 otherwise. This ensures that the model parameters are updated exclusively based on its ability to correctly predict the answer tokens.
As for the training configuration, the model was trained for a maximum of 5 epochs using the Paged AdamW optimizer, a memory-efficient variant introduced in the QLoRA framework (Dettmers et al., 2023), which builds upon the AdamW algorithm with decoupled weight decay (Loshchilov and Hutter, 2017). A constant learning rate of $2\times 10^{-5}$ was used. Performance was evaluated on the held-out validation set at the end of each epoch using the weighted F1 score as the primary metric for model selection. An early stopping mechanism with a patience of 2 was implemented to halt training if the validation F1 score did not improve for two consecutive epochs, thereby preventing overfitting and selecting the model checkpoint that best generalizes to unseen data.
@@ 3 Results and Discussion
This section presents the results of our systematic evaluation of causal LLMs in the context of transport mode choice modelling. The analysis is structured into three core components: predictive performance, reasoning quality, and finetuning effects. First, we assess the ability of eleven open access LLMs, ranging from 1B to 12B parameters, to accurately replicate human mode choice behaviour across different datasets, learning strategies, and diverse prompting and temperature configurations. Performance is evaluated at both the instance level (e.g., accuracy, macro F1) and the distributional level (e.g., Jensen-Shannon divergence), benchmarked against observed human survey data. Second, we analyze the quality and structure of the natural language rationales generated by each model. This includes both quantitative measures, such as the Explanation Strength Index (ESI), and semantic level insights extracted via topic modelling using BERTopic. The objective is to evaluate not only what the models predict, but why they make those predictions, and how closely those rationales align with human decision factors. Finally, we examine the impact of finetuning a causal LLM on labelled travel survey data. This is the first known attempt to finetune a foundation model specifically for transport mode choice prediction. We compare its performance to base models and highlight the extent to which supervised adaptation improves alignment with ground truth selections.

### v2

@@ 2.1.4 Prompting Style
To evaluate the effect of prompting style on LLM’s predictive performance, two distinct styles were implemented. The first adopts a hybrid structure inspired by Chain-of-Thought (CoT) and the Reason+Act (ReACT) prompting, wherein the model is encouraged to simulate a deliberative reasoning process before arriving at a decision. In this format, the prompt explicitly instructs the LLM to consider relevant trade-offs and contextual factors, such as time, cost, and purpose, based on the travel scenario and commuter profile, and to articulate a rationale before stating its selected mode. This approach facilitates cognitive transparency by allowing the model to express its internal reasoning prior to outputting a final decision.
@@
The second prompting style adopts a direct decision format, wherein the LLM is presented with the full scenario description including commuter attributes and available alternatives, and is asked to output only the selected mode, without generating intermediate reasoning. This method prioritizes succinctness and emulates standard classification behaviour, allowing for rapid assessments of model preference without the cognitive load in the process. Worth noting that in both styles, the models are still instructed to output a separate JSON briefly outlining the rationale behind their selected choice. The difference here lies in the intermediary process, where the first style requires the model to follow a multi-step reasoning prior their answer, while the second uses more direct approach.
2.1.5 Temperature Settings
@@
In addition to prompt style, the temperature parameter of the LLM was varied to explore the influence of generative randomness on prediction outcomes. In this context, temperature controls the stochasticity of the model’s token sampling process, with lower values (e.g., 0.5) promoting more deterministic with focused outputs, and higher values (e.g., 1.0) allowing for greater variation and exploratory generation. Temperature $0.5$ was selected to assess performance under conservative, reproducible conditions, while temperature $1.0$ was included to evaluate whether increased generative diversity may enhance generalization, particularly in cases involving ambiguity or competing trade-offs in the mode choice scenario.
@@
$11\,\text{models}\times 3\,\text{datasets}$ $\times 3\,\text{learning styles}\times 2\,\text{prompting styles}\times 2\,\text{temperatures}$
$=396\,\text{unique experiments}$
@@
$396\,\text{configurations}\times 200\,\text{test instances}=79{,}200\,\text{ Mode choice decisions}$
Each experiment was subsequently evaluated by comparing the model generated predictions against the actual survey respondents’ mode choices. This evaluation framework allows for a structured and systematic assessment of each LLM’s predictive performance under varying learning strategies, prompting styles, and temperature settings. This allows the investigation into the behavioural fidelity of LLMs in emulating human travel decision making and highlights the conditions under which these models can offer the most accurate and generalizable representations of mode choice behaviour.
@@ 2. Targeted Similarity-Based Sampling
This method selects the top-$k$ training samples most similar to agent $i$ based on structured data $d_{i}$. As defined in Section 2.1, Equation 1.
The top-$k$ training examples with the highest similarity scores are selected as:
@@
Figure 2 illustrates the full local inference setup used in this study. It presents the pipeline from structured survey data ingestion to LLM querying and storage of predictions.
Figure 2: Localized LLM and Server Setup
@@ 2.3 Performance Evaluation Framework
In this framework, LLMs are evaluated for both prediction and reasoning. The evaluation comprises two complementary components: (i) predictive performance assessment, which measures accuracy and distributional alignment with observed mode choices (Section 2.3.1), and (ii) reasoning performance assessment, which examines the structure, strength, and thematic coherence of the model-generated explanations (Section 2.3.2).
@@ 2.3.1 LLMs Predictive Performance Evaluation
Let $\mathcal{D}_{\text{test}}=\{(d_{i},y_{i})\}_{i=1}^{N}$ be the test dataset, where $d_{i}$ is the input data (trip and sociodemographic attributes), and $y_{i}\in\{1,2,\dots,C\}$ denotes the ground-truth mode label and $C$ denote the total number of transport mode categories used in the classification task. Each evaluated model $\mathcal{M}^{(m)}$ generates a predicted mode $\hat{y}_{i}^{(m)}$ for each instance $i$. We compute the following standard classification metrics:
@@
Here, $\text{TP}_{c}^{(m)}$ is the number of true positives, and $\text{FP}_{c}^{(m)}$ is the number of false positives for class $c$.
@@
$\text{FN}_{c}^{(m)}$ denotes the number of false negatives for class $c$.
$\text{F1}^{(m)}=\frac{1}{C}\sum_{c=1}^{C}\frac{2\cdot\text{Precision}_{c}^{(m)}\cdot\text{Recall}_{c}^{(m)}}{\text{Precision}_{c}^{(m)}+\text{Recall}_{c}^{(m)}}$ (17)
$\text{F1}^{(m)}_{\text{weighted}}=\sum_{c=1}^{C}w_{c}\cdot\text{F1}_{c}^{(m)}$ (18)
The weighted F1 Score accounts for class imbalance by weighting each class-specific F1 Score by its support $w_{c}=\frac{n_{c}}{N}$, where $n_{c}$ is the number of ground truth instances in class $c$. In addition to these instance-level metrics, we compute distribution-level divergence between the predicted and true mode shares. Let the empirical ground-truth distribution over classes be defined as
$p_{c}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(y_{i}=c),$ (19)
and the predicted distribution by model $\mathcal{M}^{(m)}$ as
$\hat{p}_{c}^{(m)}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(\hat{y}_{i}^{(m)}=c).$ (20)
To address zero-count issues in the computation of divergence-based metrics, we apply Laplace smoothing using a small constant $\varepsilon=10^{-9}$, yielding:
$p_{c}^{\text{smooth}}=\frac{p_{c}+\varepsilon}{1+3\varepsilon},\quad\hat{p}_{c}^{(m),\text{smooth}}=\frac{\hat{p}_{c}^{(m)}+\varepsilon}{1+3\varepsilon}.$ (21)
We then compute the following distribution-level evaluation metrics:
$\text{DistMAE}^{(m)}=\frac{1}{C}\sum_{c=1}^{C}\left|p_{c}-\hat{p}_{c}^{(m)}\right|.$ (22)
$\text{JSD}^{(m)}=\frac{1}{2}\left[\sum_{c=1}^{C}p_{c}^{\text{smooth}}\log\left(\frac{p_{c}^{\text{smooth}}}{M_{c}}\right)+\hat{p}_{c}^{(m),\text{smooth}}\log\left(\frac{\hat{p}_{c}^{(m),\text{smooth}}}{M_{c}}\right)\right],$ (23)
where $M_{c}=\frac{1}{2}\left(p_{c}^{\text{smooth}}+\hat{p}_{c}^{(m),\text{smooth}}\right)$ is the average distribution.
$\text{CE}^{(m)}=-\sum_{c=1}^{C}p_{c}^{\text{smooth}}\log\left(\hat{p}_{c}^{(m),\text{smooth}}\right).$ (24)
Each metric in the evaluation framework was selected for its ability to capture distinct yet complementary aspects of model performance. The accuracy metric. provides an overall correctness measure but can be biased by class imbalance. The precision and recall help evaluate model performance on individual classes, with F1_Macro offering a class-averaged harmonic mean and F1_Weighted accounting for class imbalance by weighting each class’s F1 score by its support. On the distribution level, MAE quantifies the average deviation between predicted and true mode share proportions, while cross-entropy penalizes probabilistic mismatch and is sensitive to over or under-confidence in predictions. JSD, on the other hand a symmetrized and smoothed variant of Kullback-Leibler divergence, offers a bounded and interpretable measure of similarity between the predicted and actual distributions (Lin, 2002).
@@ 2.3.2 LLMs Reasoning Performance Evaluation
To complement the evaluation of predicted travel modes, we conducted a detailed reasoning analysis to assess the structure, diversity, and strength of the natural language justifications generated by each model. This analysis consisted of two main components: (i) computation of an explanation strength index, and (ii) topic modelling using BERTopic (Grootendorst, 2022).
For each generated reasoning $z_{i}$ from instance $i$, we defined an Explanation Strength Index $\mathrm{ESI}(z_{i})$ that quantifies the extent to which the explanation references utility-relevant decision factors. Let $\mathcal{F}=\{\text{time},\text{cost},\text{comfort},\text{convenience},$ $\text{frequency}\}$ denote the set of predefined decision-related keywords. The $\mathrm{ESI}(z_{i})$ is computed as:
@@
Additionally, each reasoning $z_{i}$ was encoded into a high-dimensional semantic vector $\mathbf{e}_{i}\in\mathbb{R}^{d}$ using the all-MiniLM-L6-v2 sentence transformer model. These embeddings were reduced to two dimensions via Uniform Manifold Approximation and Projection (UMAP) for visualization:
@@
with UMAP parameters set to $n_{\text{neighbors}}=15$, $\text{min\_dist}=0.1$, and cosine distance as the metric.
The BERTopic algorithm was then applied using the sentence embeddings and UMAP-reduced vectors. This model clustered the explanations into coherent topics $T=\{t_{1},t_{2},\dots,t_{K}\}$, with each instance assigned a topic label $t_{i}\in T$. For each topic, a ranked list of representative keywords was extracted using class-based TF-IDF, for enhanced interpretability.
These different evaluation dimensions enabled a structured evaluation of the interpretive dimension of model outputs, which allowed us to comprehensively assess the content richness and diversity of the generated rationales along with the models predictive accuracy.
@@ 2.4 Fine-tunning
The foundation for this research is the decoder-only transformer LLMs a.k.a. Casual LLMs. Its extensive pre-training provides a rich basis of general knowledge, which we specialize for travel mode choice prediction task through finetuning.
The training corpus consists of several thousand structured examples derived from the Swissmetro travel survey responses except for a hold sample that was used for testing. Each example is composed of: 1) an instruction, a detailed text prompt describing a traveller’s sociodemographic profile and a specific travel scenario with the transport options and their attributes (e.g., travel time, cost, frequency); and 2) a selected_mode, the ground-truth label corresponding to the mode choice.
For robust model development, the dataset was partitioned into a training set (90%) and a validation set (10%). This allows for performance monitoring on unseen data to prevent overfitting and to select the best-performing model checkpoint. This is separate from the 200 test set that was set aside for the finetuned model validation and comparison, similar to the process described in Section 2.3.
@@ 2.4.1 Parameter Efficient Finetuning (PEFT) Strategy: QLoRA
Our parameter-efficient finetuning strategy is based on Quantized Low-Rank Adaptation (QLoRA) (Dettmers et al., 2023), an advanced method that enhances the original Low-Rank Adaptation (LoRA) framework (Hu et al., 2022) by integrating 4-bit quantization of the base model’s weights, thereby significantly reducing memory requirements. This strategy makes finetuning large models feasible on limited hardware (a single GPU with 12GB VRAM used for this research) by drastically reducing the number of trainable parameters.
@@
For this study, we configured LoRA with a rank $r=32$ and a scaling factor $\alpha=64$. These trainable matrices were applied to all major linear layers of the model architecture, including query, key, value, and output projections, as well as the feed-forward network layers.
@@ 2.4.2 Training Objective and Label Leakage Mitigation
As the model was finetuned using a Causal LLM objective, a significant challenge when applying this objective to structured prediction tasks with decoder-only architectures is label leakage (Raffel et al., 2020; Li et al., 2023). This occurs when the ground-truth answer is exposed to the model as part of its input context, allowing the model to achieve high performance by simply copying the answer rather than learning the task.
To resolve this, we implemented a loss masking strategy. While the model receives the full sequence $\text{Prompt},\text{Answer}]$ as input to maintain conversational context, the loss function is computed only for the Answer portion. Let a sequence of tokens be $\boldsymbol{\tau}=(\tau_{1},\tau_{2},\dots,\tau_{|Q|})$ and its corresponding labels be $\boldsymbol{\lambda}=(\lambda_{1},\lambda_{2},\dots,\lambda_{|Q|})$, where the labels for prompt tokens are set to an ignore index of 100-100. The masked cross-entropy loss is formulated as:
@@
where $\theta$ represents the model’s trainable parameters, $P(\tau_{j}\mid\tau_{<j};\theta)$ is the predicted probability of token $\tau_{j}$, and $\mathbb{I}(\cdot)$ is an indicator function that equals 1 if the condition is true and 0 otherwise. This ensures that the model parameters are updated exclusively based on their ability to correctly predict the answer tokens.
As for the training configuration, the model was trained for a maximum of 5 epochs using the Paged AdamW optimizer, a memory-efficient variant introduced in the QLoRA framework (Dettmers et al., 2023), which builds upon the AdamW algorithm with decoupled weight decay (Loshchilov and Hutter, 2017). A constant learning rate of $2\times 10^{-5}$ was used. Performance was evaluated on the held-out validation set at the end of each epoch using the weighted F1 score as the primary metric for model selection. An early stopping mechanism with a patience of 2 was implemented to halt training if the validation F1 score did not improve for two consecutive epochs to prevent overfitting and to select the model checkpoint that best generalizes to unseen data. Worth noting, that the reported F1 and JSD scores for the fine-tuned model in section 3.4 are based on the model deployment on the test data, which were not part of the training or the validation data used at the fine-tuning stage.
@@ 3 Results and Discussion
This section presents the results of our systematic evaluation of causal LLMs in the context of transport mode choice modelling. The analysis is structured into three core components: predictive performance, reasoning quality, and finetuning effects. First, we assess the ability of eleven open access LLMs, ranging from 1 to 12B parameters, to accurately replicate human mode choice behaviour across different datasets, learning strategies, and diverse prompting and temperature configurations. Performance is evaluated at both the instance level (e.g., accuracy, macro F1) and the distributional level (e.g., Jensen-Shannon divergence), benchmarked against observed human survey data. Second, we analyze the quality and structure of the natural language rationales generated by each model. This includes both quantitative measures, such as the Explanation Strength Index (ESI), and semantic level insights extracted via topic modelling using BERTopic. The objective is to evaluate not only what the models predict, but why they make those predictions, and how closely those rationales align with human decision factors. Finally, we examine the impact of fine-tuning a causal LLM on labelled travel survey data. This is the first known attempt to fine-tune a foundation model specifically for transport mode choice prediction. We compare its performance to base models and highlight the extent to which supervised adaptation improves alignment with ground truth selections.